OpenStudy (anonymous):

Find the limit as: lim (x-3)/(3x-1) x->00 I am having a major brian fart here and wanted to know if I solved this right.

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}\left(\begin{matrix}(x-1) \\ (3x-1)\end{matrix}\right)\]

OpenStudy (anonymous):

Its suppose to be a fraction

OpenStudy (amistre64):

the limit of a rational funtion as x increase is the horisontal asymptote if it can be found

OpenStudy (anonymous):

ok so this one would be 1/3

OpenStudy (amistre64):

since the bottom is the same degree as the top; its the ration of leading coeffs

OpenStudy (amistre64):

1/3 is correct

OpenStudy (anonymous):

ok thank you very much

OpenStudy (amistre64):

in essense you get for large xs: \[\frac{1(10000000)}{3(10000000)}\]

OpenStudy (anonymous):

Yea I know we don't worry about the others because they essentially 0 out, well according to my professor anyways. I had a calc test early today I could not for the life of me think of the correct way to solve, I knew how I just could not do it; it was very frustrating. But thank you for your help